OpenStudy (anonymous):

in triangle ABC, controid D is on median line AM. AD=x+5 and DM=3x-5. what is AM?

OpenStudy (anonymous):

Can you draw the triangle ?

OpenStudy (anonymous):

uhhh..i dont know how to thank you anyway...

OpenStudy (anonymous):

Wait i am drawing it i solved it i guess

OpenStudy (anonymous):

 \[3k=x+5\] --> AD \[k=(x+5)/3\] \[k=3x+5\] --> DM \[3(3x-5)=x+5\] \[6x-15=x+5\] \[5x=20\] \[x=4\]